OpenStudy (anonymous):

Michelle is creating a number pattern. She begins with the number 2. Each number after 2 is three times the number before it, minus 2. Which pattern fits Michelle’s rule? a. 2, 4, 10, 28, 82 b. 2, 6, 10, 54, 82 c. 2, 4, 18, 28, 162 d. 2, 6, 18, 54, 162

OpenStudy (anonymous):

a. because 2*3=6-2=4 4*3=12-2=10 etc.

OpenStudy (anonymous):

thx